# Distributed Learning Dynamics Converging to the Core of -Matchings

### v1

@@ VI-A Sample run and average performance
The illustrations shown are algorithm runs with a setup of the following main parameters:
β’
tasks with ranging from to
β’
robots with ranging from to
β’
The left side of Fig. 1 shows the final matching between the robots nodes, with each node expanded as copies, and the tasks nodes, with each node expanded as copies. The right side of Fig. 1 has a plot of the total aspirations of all the copies involved in a matching at each iteration. The dashed line represents the maximum -matching value solved by a linear program.
In Fig. 2, we ran the algorithm for different configurations, with the same parameters as above but with randomization of the values and -values of the tasks. The figure shows the plot of the average of the relative total feasible aspirations, where, for each run, the relative total feasible aspiration is the total feasible aspirations divided by the optimal value for that run.
@@ VI-B Parameters effects
We ran various simulations to help us gain insights into the effects of some of the problem parameters on the average performance of the dynamics, specifically the -values, number of nodes, and discretization size. Figures 3, 4, and 5 illustrate the observed trends resulting from the parameters variations.
First, to isolate the effect of varying the -value, we generated random -matching instances, and for each instance, we ran the dynamics for different -value functions, and . value is randomly generated such that is such that and is such that . Fig. 3 shows the average of the relative welfare across runs while using the different -value functions. Empirically, it appears for this setup that bigger -values lead to better performance of the dynamics, on average.
@@ VII CONCLUSION
To recap, we considered the -matching setup. We defined a copies core solution, which is a solution concept that specifies the gain of each node in a bipartite graph, from each match it joins. The copies-core solution achieves core stability such that no group of agents can deviate and collectively gain a higher payoff through another matching. We presented a centralized algorithm, in the nature of the Hungarian method, that reaches the copies-core solution in polynomial time. We further introduced distributed learning dynamics that converge to the core with probability one. Finally, we illustrated the distributed dynamics on a multi-robot task assignment setting.

### v2

@@ VI-A Sample run and average performance
Fig. 1 illustrates a sample run of the dynamics on the task assignment setting. The left side of Fig. 1 shows the final matching between the robot nodes, with each node expanded as copies, and the task nodes, with each node expanded as copies. The right side of Fig. 1 is a plot of the total feasible aspirations, aspirations of the matched copies, at each iteration. The dashed line represents the maximum -matching value solved by a linear program.
To observe the average performance of the dynamics on the preceding setting, we ran the dynamics for different configurations, setting and randomizing the and -values of the tasks. Fig. 2 shows the plot of the average of the relative total feasible aspirations, where, for each run, the relative total feasible aspiration is the total feasible aspirations divided by the optimal value for that run.
@@ VI-B Parameters effects
We ran supplementary simulations for general -matching problems to help us gain insights into the effects of some of the problem parameters on the average performance of the dynamics, specifically the -values, number of nodes, and discretization size. Figures 3, 4, and 5 illustrate the observed trends resulting from the parameters variations.
First, to isolate the effect of varying the -values, we generated random -matching instances; we set the number of nodes per each class to be yet varied the range of -values. For each instance, we ran the dynamics for different -value functions, and . value is randomly generated such that is such that and is such that . Fig. 3 shows the average of the relative welfare across the runs while using the different -value functions. Empirically, it appears for this setup that, on average, bigger -values lead to better performance of the dynamics.
@@ VII CONCLUSION
To recap, we considered the -matching setup. We defined a copies-core solution, which is a solution concept that specifies the gain of each node in a bipartite graph, from each match it joins. The copies-core solution achieves core stability such that no group of agents can deviate and collectively gain a higher payoff through another matching. We presented a centralized algorithm, in the nature of the Hungarian method, that reaches the copies-core solution in polynomial time. We further introduced distributed learning dynamics that converge to the core with probability one. Finally, we illustrated the distributed dynamics on a multi-agent task assignment setting.